OpenStudy (anonymous):

subtract ? 4√5 - 9√20

OpenStudy (anonymous):

You'll want to simplify \(\sqrt{20}\) before you try to add.

OpenStudy (anonymous):

\[-14 \sqrt{5} \]\[4\surd 5-9\surd 20 \]\[4\surd 5-18 \sqrt{5} \]\[-14 \sqrt{5} \]

OpenStudy (anonymous):

what ? to both of your responses. pol im subtracting. and robt whered you get 14 ?

OpenStudy (anonymous):

4 apples take away 18 apples leaves negative 14 apples.

OpenStudy (anonymous):

whered 18 come from?

OpenStudy (anonymous):

Simplifying \(9\sqrt{20}\) like I said: \[9\sqrt{20} = 9 \sqrt{4\cdot 5} = 9 \cdot\sqrt{4} \cdot \sqrt{5} = 9\cdot 2 \cdot \sqrt{5} = 18 \sqrt{5}\]

OpenStudy (anonymous):

\[9\sqrt{20}=9\sqrt{2*2*5}=9*2\sqrt{5}=(18)\sqrt{5} \]I was working in parallel with polpak.